OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11

OpenStudy (anonymous):

You want the coefficients on x to be opposites. What would you multiply by the first equation so the coefficient on x is the opposite of that on the second equation?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@JackJones -3? 0r 3?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

6 + (-6) = 0

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

Idk?? You think you could answer a few more?

OpenStudy (anonymous):

if i know it

OpenStudy (anonymous):

go ahead and post another question, but first give me a medal and close this one

OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6 3 2 6 none of the above

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Solve the system using the addition method. 3x + 4y = -10 5x - 2y = 18 What is the x-value in the solution? 2 -2 -4 none of the above

OpenStudy (anonymous):

Solve the system using the addition method. 3x + 2y = 10 6x + 4y = 14 What is the solution to the system? (2, 2) (1, 2) Infinite solutions on the line No solution

OpenStudy (anonymous):

After those two I'm done.

OpenStudy (anonymous):

2 and no solution

OpenStudy (anonymous):

Ahh.. so sorry there's one more.

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

Solve the system using the addition method. 5x - 2y = 3 2x - y = 0 What is the y-value in the solution? 3 6 -3 none of the above

OpenStudy (anonymous):

6

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

Thank you!!